OpenStudy (anonymous):

3. A 3.6 µC point charge is moved within an electric field and has an electric potential energy change of 14.2 J. What is the electric potential difference before and after the charge was moved?

OpenStudy (anonymous):

I get that V*Q= Energy so V must equal Energy/Q

OpenStudy (michele_laino):

here we have to apply this formula: \[\large \Delta V = \frac{{\Delta U}}{q} = \frac{{14.2}}{{3.6 \times {{10}^{ - 6}}}}\]